OpenStudy (anonymous):

What is the integral of: (1+(cosx)^2)^0.5

OpenStudy (anonymous):

\[\int \sqrt{1+\cos^2 x} \ \text{d}x\]???

OpenStudy (anonymous):

if yes there is no closed form for that integral in terms of elementary functions

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

between 0 and pi